# Networking Systems for Video Anomaly Detection: A Tutorial and Survey

### v2

@@ 1.4. Section Navigation
Based on the relationship between the content of the individual sections and the NSVAD architecture shown in Fig. 1, we organize the remainder of this article and provide an intuitive navigation map shown in Fig. 3. Specifically, Section 2 states the general basics of VAD, including the task definition, type of anomalies, and application areas. Sections  3$\sim$5 elaborate on the learning paradigms and typical models of UVAD, WsVAD and FuVAD, respectively. We provide detailed explanations of classic methods for readers to understand the core ideas and implementations better. To assist readers to conduct research instantly, we provide a comprehensive introduction to existing datasets and evaluation metrics involved in the current work in Section 6. Section 7 presents our latest exploration of NSVAD, including how to develop high-performance VAD models for dynamic environments in industrial IoT and how to address arbitrary data bias in smart city scenarios. Finally, Section 8 discusses the future vision of NSVAD, providing an in-depth analysis of its existing challenges, development trends, and possible opportunities. This tutorial paper is intended for non-specialists, so we prioritize conceptual clarification and research horizon construction in the main papers. Content that may overlap with existing work (e.g., itemized introductions to reviewed papers, detailed explanations of classical methods, and comprehensive presentations of research cases) is relegated to the supplement.
@@ 2. Generalized Foundations of NSVAD
As a cross-cutting topic, NSVAD has attracted researchers from deep learning, video surveillance, mobile internet, and edge computing communities. Initially, VAD followed the conventional setting of AD problems, where anomalies were treated as outliers with different distributions (Ramachandra et al., 2020). Corresponding unsupervised methods (Cao et al., 2024; Han et al., 2024b; Ye et al., 2024) aimed to learn prototypical representations of regular events, considering test videos outside the distribution as anomalies. To address video’s high-dimensional and complex backgrounds, AD researchers introduced efficient video representation learning techniques like Auto-Encoders (AEs) (Park et al., 2020), Generative Adversarial Networks (GANs) (Sun et al., 2024; Huang et al., 2023; Wu et al., 2024a), Transformers (Feng et al., 2021b), and diffusion models (Basak and Gautam, 2024). With large-scale video datasets and high-performance GPUs, deep learning-driven Video Understanding (VU) techniques have advanced, shifting VAD research to cross-cutting topics of VU and AD. New VAD routes emerged, such as WsVAD with multiple instance learning (Sultani et al., 2018) and FuVAD with iterative learning (Pang et al., 2020). These methods challenge the open-world assumption under unsupervised AD, where real-world anomalies are varied and unbounded. WsVAD incorporates anomaly instances in training to differentiate between negative and positive samples. While WsVAD requires collecting and labeling anomalies, it outputs more reliable results for specific types of abnormal events. FuVAD avoids the data constraints of UVAD and WsVAD by learning anomaly detectors directly from raw videos, reducing data preparation costs and preventing mislabeling issues.
In recent years, the rise of multimodal learning (Xu et al., 2023; Ektefaie et al., 2023) and large language models (Radford et al., 2021a; Reddy et al., 2021) have brought a new windfall for VAD. Researchers have proposed new tasks such as open-set VAD (Acsintoae et al., 2022; Zhu et al., 2022), open-vocabulary VAD (Wu et al., 2024c), and Video Anomaly Retrieval (VAR) (Wu et al., 2024b), indicating the trend of integration between VAD and generative AI research. We understand the cognitive differences resulting from the research backgrounds of these fields and view the emergence of new routes and tasks as a positive signal to promote VAD to systematic NSVAD research. In addition, NSVAD systems deployed in real-world scenarios must face the challenge of domain bias due to multi-view, cross-scenario videos and consider the limited storage and communication resources of end devices. Recent NSVAD advances (Liu et al., 2024b; Wu et al., 2020) have begun considering both algorithm optimization and model deployment to balance performance and overhead.
@@ 2.1. Related Terms
We introduce key AD-related terms, including Anomaly Detection (AD), Novelty Detection (ND), Open Set Recognition (OSR), and Outlier Detection (OD). These terms often confuse researchers from the computer vision community (Wang et al., 2024a; Liu et al., 2024c), and AD practitioners struggle with inconsistent definitions (Yang et al., 2021). We follow community consensus and our experience to clarify these terms, as shown in Fig. 4.
AD detects samples that deviate from normality as defined by training data (Samariya and Thakkar, 2023). Such deviations include: 1) covariate shift, i.e., label-independent distributional differences due to factors like image style or equipment, and 2) semantic shift, i.e., samples from different categories. The former is addressed by Sensory Anomaly Detection (SenAD), which includes tasks like domain adaptation. UVAD, by contrast, focuses on detecting anomalous samples with different semantic labels, termed Semantic AD (SemAD).
@@ 2.3. Definition of Various NSVAD Routes
UVAD refers to NSVAD schemes that use only easily collected routine events to train models to learn the spatial-temporal pattern boundaries of normal samples (Ramachandra et al., 2020; Zhang et al., 2024a). UVAD dominated early VAD research because it follows the open-world assumptions in the same vein as the AD community, circumventing predefinitions and collecting anomalous instances (Liu et al., 2021a; Huang et al., 2024b). In Fig. 5, we show the general learning framework of UVAD and compare it with WsVAD and FuVAD. Specifically, UVAD assumes that models trained on regular events will only describe normal spatial-temporal patterns and will exhibit significant deviations when confronted with unseen anomalous examples, such as probability distributions (Sabokrou et al., 2015; Saligrama et al., 2010; Cheng et al., 2015; Benezeth et al., 2009), distances (Dalal et al., 2006; Saligrama and Chen, 2012; Dalal and Triggs, 2005), and proxy task errors (Liu et al., 2022c; Hasan et al., 2016; Gong et al., 2019; Liu et al., 2024b). Early approaches first used local binary operators (Hu et al., 2018; Nawarathna et al., 2014; Zhang et al., 2015), spatial-temporal points of interest (Dollár et al., 2005), etc. to characterize spatial-temporal features that are normal events, and then employed One-Class (OC) classifiers (e.g., OC support vector machines and OC neural networks) (Ionescu et al., 2019; Sun et al., 2019) to learn the pattern boundaries, and considered test samples whose features fell outside the boundaries as anomalous. Such methods rely on manual features and are prone to dimensionality disasters.
@@
In recent years, deep learning-driven UVAD integrates feature extraction and normality learning into a unified framework with two phases, training and testing, which corresponds to normality learning via the use of negative samples and anomaly detection via the detection of out-of-distribution samples (Liu et al., 2024d). In the normality learning phase, the network learnable parameters are optimized by minimizing a loss function overall negative samples. Whereas, in the testing phase, the degree of anomaly is measured by quantifying the distance between the test samples and the learned normality. Among them, reconstruction-based methods have dominated UVAD research in recent years (Liu et al., 2018b; Gong et al., 2019; Park et al., 2020; Liu et al., 2022c, 2024b). On the one hand, most of the challenges faced by such methods, such as global motion modeling, temporal normality learning, and spatial detail inference, have been intensively studied in video self-supervised learning. As a result, many methods have driven the development of UVAD by drawing inspiration from existing methods, such as video prediction (Oprea et al., 2020; Wang et al., 2022). On the other hand, since reconstruction/prediction methods aim to learn a generative model that can reason about regular events, their basic settings and optimization goals are clear and unambiguous, and thus easy to implement and follow. Essentially, UVAD is transductive learning, i.e., model training and testing are relatively independent. Numerous studies have shown that the performance of models in the normality learning phase on the agent task does not show a positive correlation with downstream anomaly detection. Due to the diversity of events, the spatial-temporal features of normal and abnormal samples overlap, and UVAD usually fails to actively recognize discriminative features and incorrectly learns some shared patterns when only regular events are available for model training (Liu et al., 2023e). In addition, Park et al. (Park et al., 2020) pointed out that overpowered deep neural networks may be able to effectively reason about unseen anomalous events during the testing phase due to overgeneralization performance, which may lead to underdetection. In response, the researchers proposed a memory network enhancement approach to weaken the model’s ability to generalize representations of anomalies by recording prototypical features.
@@
WsVAD uses weakly semantic video-level labels to supervise the output of strong semantic frame-level labels, i.e., frame-by-frame anomaly scores, by the sequencing model, thus enabling temporal localization of anomalous events (Liu et al., 2022b; Karim et al., 2024). The first weakly-supervised approach is the multiple instance ranking framework proposed by Sultani et al. (Sultani et al., 2018) in 2018, which lays out the basic MIL architecture of the WsVAD route, and whose concurrently publicly available UCF-Crime dataset has become the most widely used weakly-supervised benchmark. They consider a video as a collection of multiple examples (video clips), where clips containing abnormal frames are positive examples, while clips with all normal frames are labeled as negative examples. Obviously, a normal video with label 0 produces all examples called negative bags, and the example-level labels are all 00. An abnormal video with the label $1$ constitutes a positive packet, which contains both positive and negative instances. Inspired by multiple instance learning, WsVAD aims to train a scoring model to output the anomaly scores of each example using video-level labels. The authors introduce a MIL ranking loss inspired by the hinge loss, which encourages the model to output high anomaly scores close to 1 for anomalous clips by maximizing the difference between the anomaly scores of the largest-scoring instances in the positive and negative bags while scoring regular clips as close to 0 as possible. In fact, WsVAD does not belong to any of the classes of out-of-distribution detection tasks introduced in Section 2.1, but is rather a type of multiple instance learning under weak semantic labeling supervision.
@@
Compared to UVAD, the weakly supervised approach introduces anomalous videos in the training set and provides video-level labels for all training samples. Although anomalous examples are diverse and unenumerable in the real world, noteworthy anomalous events in specific scenarios are usually limited and easy to collect, such as thefts, robberies, and traffic accidents, which can be obtained in large quantities from surveillance IoT systems and online video platforms. In addition, the labor cost of video-level tagging, which only requires marking whether a video contains an anomaly without worrying about the specific timing location (frame-level labeling) and detailed anomaly category, is usually affordable. For example, the UCF-Crime dataset is much larger than the one used by UVAD, but only 1,900 discrete $[0,1]$ labels need to be provided. Due to the introduction of an additional human prior and the fact that the model has seen anomalous events during the training phase, WsVAD results are typically more reliable than UVAD, achieving excellent and consistent performance in detecting specific anomalous events. As a result, WsVAD has become a mainstream VAD scheme and is considered as the most promising research route for deployment in intelligent surveillance systems. The latest research attempts to mine anomaly-related clues from audio or subtitle text accompanying video frames, and proposes the multimodal WsVAD (Yang et al., 2024b). This article introduces the recent progress of WsVAD in Section 4.
@@
FuVAD attempts to learn anomaly classifiers directly from large-scale raw videos without any editing and labeling (Pang et al., 2020; Zaheer et al., 2022). Specifically, FuVAD’s training data contains both positive and negative samples, and due to the low frequency of anomalous examples compared to regular events, the anomalous samples to be detected can be regarded as outliers with different patterns from the main data. In essence, the FuVAD model is transductive learning and does not follow the training-testing process of UVAD. Fully unsupervised methods have become a research hotspot in the internet era, where data preparation is costly, by virtue of the fact that they do not require any constraints and assumptions on the training data, and can be used to train the model by directly accessing huge amounts of videos from the real world (Liu et al., 2024d).
@@ 2.4. Emerging Research Tasks
2.5. Language model-based VAD
Language models like Generative Pre-Training (GPT) (Radford, 2018) and Contrastive Language-Image Pretraining (CLIP) (Radford et al., 2021b) exhibit outstanding zero-shot learning and multimodal information processing abilities, showing great potential in VAD. Language models (Yang et al., 2024c; Teng et al., 2024) can provide accurate textual descriptions of video frames, enhancing the semantic mining capability of existing vision-based VAD models. For instance, Chen et al. (Chen et al., 2023) used a language model-based captioning network to obtain textual descriptions of video sequences, which, after being embedded in a text embedding network, were fused with visual features as inputs to the anomaly detector. Their proposed Text Empowered Video Anomaly Detection (TEVAD) efficiently captures abstract semantics of anomaly events and enhances the interpretability of VAD models. Pu et al. (Pu et al., 2024) introduced a Prompt-Enhanced Learning (PEL) module, using knowledge-based prompts to incorporate semantic priors, improving the discriminative power of visual features in weakly supervised VAD, while ensuring separability between anomalous subcategories. Wu et al. (Wu et al., 2024b) proposed VAR, which efficiently detects specific anomalies based on cross-modal learning (e.g., language descriptions and synchronized audio). They designed an Anomaly-Led Alignment Network (ALAN) using BERT (Kenton and Toutanova, 2019) to process text information and incorporate a pretext task to enhance semantic alignment between video-text fine-grained representations.
In addition, visual-language associations can serve as effective cues for detecting video anomalies, and the pre-learned visual-text consistency in large Visual-Language Models (VLMs) can be efficiently transferred to VAD. Kim et al. (Kim et al., 2023) used large language models to generate textual descriptions of video frames and detected anomalous frames by calculating the cosine similarity between input frames and their textual descriptions using CLIP. Text Prompt with Normality Guidance (Yang et al., 2024b) leverages the language-visual knowledge of the CLIP model to align video frames with textual descriptions of events, generating more accurate pseudo-labels for WsVAD, thus improving model performance. Inspired by CLIP, Wu et al. (Wu et al., 2024e) proposed a dual-branch structure called VadCLIP that utilizes the connection between vision and language for weakly supervised VAD. One branch uses only visual features for coarse-grained binary classification, while the other fully exploits fine-grained visual-language alignment to transfer CLIP’s knowledge to VAD. They also used a pre-trained VLM to learn spatio-temporal prompt embeddings for weakly supervised VAD and localization. The proposed STPrompt (Wu et al., 2024d) leverages the knowledge learned from VLMs and integrates natural motion priors from the original video to learn prompt embeddings consistent with spatial-temporal regions of the video (e.g., patches of individual frames) to identify specific local regions of anomalies.
@@ 2.5. Language model-based VAD
Research has shown that multimodal LLMs can learn prototype patterns of normal events in a zero-shot manner and describe any anomalies in open-set settings, significantly improving the generality and adaptability of VAD models. Zanella et al. (Zanella et al., 2024) proposed Language-based VAD (LAVAD), an unsupervised learning paradigm that leverages pre-trained LLMs and existing VLMs to train video anomaly detectors. They used VLMs to generate textual descriptions of video frames and designed a prompt mechanism to unlock LLMs’ potential in temporal aggregation and anomaly score estimation, enabling direct VAD execution. Lv et al. (Lv and Sun, 2024) introduced Video-LLaMA into VAD, aiming to break threshold limitations and improve model interpretability. The authors proposed a three-stage training method to improve VLLM’s training efficiency. In the AnomalyRuler, during the inductive stage, a small set of normal reference videos was provided to the LLM, enabling it to summarize normal patterns to induce rules for anomaly detection. In the Deduction stage, these induced rules were applied to detect anomalous frames in test videos. Hawk in (Tang et al., 2024) uses an interactive VLM to accurately interpret video anomalies, answering VAD-related questions. It constructs an auxiliary consistency loss within the motion and video space, guiding the video branch to focus on motion modalities and establishing explicit supervision between actions and language to improve the accuracy of interpretation. Zhang et al. (Zhang et al., 2024c) developed a large-scale multimodal VAD instruction tuning benchmark called VAD-Instruct50k, used to build unbiased and interpretable VAD systems, as well as Holmes-VAD for anomaly event localization and interpretation.

### v3

@@ 1.4. Section Navigation
Based on the relationship between the content of the individual sections and the NSVAD architecture shown in Fig. 1, we organize the remainder of this article and provide an intuitive navigation map shown in Fig. 3. Specifically, Section 2 states the general basics of VAD, including the task definition, type of anomalies, and application areas. Sections  3$\sim$5 elaborate on the learning paradigms and typical models of UVAD, WsVAD and FuVAD, respectively. We provide detailed explanations of classic methods for readers to understand the core ideas and implementations better. To assist readers to conduct research instantly, we provide a comprehensive introduction to existing datasets and evaluation metrics involved in the current work in Section 6. Finally, Section 7 discusses the future vision of NSVAD, providing an in-depth analysis of its existing challenges, development trends, and possible opportunities. This tutorial paper is intended for non-specialists, so we prioritize conceptual clarification and research horizon construction in the main papers. Content that may overlap with existing work (e.g., itemized introductions to reviewed papers, detailed explanations of classical methods, and comprehensive presentations of research cases) is relegated to the Supplement.
@@ 2. Generalized Foundations of NSVAD
As a cross-cutting topic, NSVAD has attracted researchers from deep learning, video surveillance, mobile internet, and edge computing communities. Initially, VAD followed the conventional setting of AD problems, where anomalies were treated as outliers with different distributions (Ramachandra et al., 2020). Corresponding unsupervised methods (Cao et al., 2024a; Han et al., 2024b; Ye et al., 2024) aimed to learn prototypical representations of regular events, considering test videos outside the distribution as anomalies. To address video’s high-dimensional and complex backgrounds, AD researchers introduced efficient video representation learning techniques like Auto-Encoders (AEs) (Park et al., 2020), Generative Adversarial Networks (GANs) (Sun et al., 2024; Huang et al., 2023; Wu et al., 2024a), Transformers (Feng et al., 2021b), Mamba (Li et al., 2024c), and diffusion models (Basak and Gautam, 2024; Liu et al., 2025a; Li et al., 2023). With large-scale video datasets and high-performance GPUs, deep learning-driven Video Understanding (VU) techniques have advanced, shifting VAD research to cross-cutting topics of VU and AD. New VAD routes emerged, such as WsVAD with multiple instance learning (Sultani et al., 2018) and FuVAD with iterative learning (Pang et al., 2020). These methods challenge the open-world assumption under unsupervised AD, where real-world anomalies are varied and unbounded. WsVAD incorporates anomaly instances in training to differentiate between negative and positive samples. While WsVAD requires collecting and labeling anomalies, it outputs more reliable results for specific types of abnormal events. FuVAD avoids the data constraints of UVAD and WsVAD by learning anomaly detectors directly from raw videos, reducing data preparation costs and preventing mislabeling issues.
In recent years, the rise of multimodal learning (Xu et al., 2023; Ektefaie et al., 2023) and large language models (Radford et al., 2021) have brought a new windfall for VAD. Researchers have proposed new tasks such as open-set VAD (Acsintoae et al., 2022; Zhu et al., 2022), open-vocabulary VAD (Wu et al., 2024c), and Video Anomaly Retrieval (VAR) (Wu et al., 2024b), indicating the trend of integration between VAD and generative AI research. We understand the cognitive differences resulting from the research backgrounds of these fields and view the emergence of new routes and tasks as a positive signal to promote VAD to systematic NSVAD research. In addition, NSVAD systems deployed in real-world scenarios must face the challenge of domain bias due to multi-view, cross-scenario videos and consider the limited storage and communication resources of end devices. Recent NSVAD advances (Liu et al., 2024d; Wu et al., 2020) have begun considering both algorithm optimization and model deployment to balance performance and overhead.
@@ 2.1. Related Terms
We introduce key AD-related terms, including Anomaly Detection (AD), Novelty Detection (ND), Open Set Recognition (OSR), and Outlier Detection (OD). These terms often confuse researchers from the computer vision community (Wang et al., 2024a; Liu et al., 2024e), and AD practitioners struggle with inconsistent definitions (Yang et al., 2021). We follow community consensus and our experience to clarify these terms, as shown in Fig. 4.
AD detects samples that deviate from normality as defined by training data (Liu et al., 2022c). Such deviations include: 1) covariate shift, i.e., label-independent distributional differences due to factors like image style or equipment, and 2) semantic shift, i.e., samples from different categories. The former is addressed by Sensory Anomaly Detection (SenAD), which includes tasks like domain adaptation. UVAD, by contrast, focuses on detecting anomalous samples with different semantic labels, termed Semantic AD (SemAD).
@@ 2.3. Definition of Various NSVAD Routes
UVAD refers to NSVAD schemes that use only easily collected routine events to train models to learn the spatial-temporal pattern boundaries of normal samples (Ramachandra et al., 2020; Zhang et al., 2024a). UVAD dominated early VAD research because it follows the open-world assumptions in the same vein as the AD community, circumventing predefinitions and collecting anomalous instances (Liu et al., 2021a; Huang et al., 2024b). In Fig. 5, we show the general learning framework of UVAD and compare it with WsVAD and FuVAD. Specifically, UVAD assumes that models trained on regular events will only describe normal spatial-temporal patterns and will exhibit significant deviations when confronted with unseen anomalous examples, such as probability distributions (Sabokrou et al., 2015; Saligrama et al., 2010; Cheng et al., 2015; Benezeth et al., 2009), distances (Dalal et al., 2006; Saligrama and Chen, 2012; Dalal and Triggs, 2005), and proxy task errors (Liu et al., 2022e; Hasan et al., 2016; Gong et al., 2019; Liu et al., 2024d). Early approaches first used local binary operators (Hu et al., 2018; Nawarathna et al., 2014; Zhang et al., 2015), spatial-temporal points of interest (Dollár et al., 2005), etc. to characterize spatial-temporal features that are normal events, and then employed One-Class (OC) classifiers (e.g., OC support vector machines and OC neural networks) (Ionescu et al., 2019; Sun et al., 2019) to learn the pattern boundaries, and considered test samples whose features fell outside the boundaries as anomalous. Such methods rely on manual features and are prone to dimensionality disasters.
@@
In recent years, deep learning-driven UVAD integrates feature extraction and normality learning into a unified framework with two phases, training and testing, which corresponds to normality learning via the use of negative samples and anomaly detection via the detection of out-of-distribution samples (Liu et al., 2024h). In the normality learning phase, the network learnable parameters are optimized by minimizing a loss function overall negative samples. Whereas, in the testing phase, the degree of anomaly is measured by quantifying the distance between the test samples and the learned normality. Among them, reconstruction-based methods have dominated UVAD research in recent years (Liu et al., 2018b; Gong et al., 2019; Park et al., 2020; Liu et al., 2022e, 2024d). On the one hand, most of the challenges faced by such methods, such as global motion modeling, temporal normality learning, and spatial detail inference, have been intensively studied in video self-supervised learning. As a result, many methods have driven the development of UVAD by drawing inspiration from existing methods, such as video prediction (Oprea et al., 2020; Wang et al., 2022b; Zhao et al., 2025). On the other hand, since reconstruction/prediction methods aim to learn a generative model that can reason about regular events, their basic settings and optimization goals are clear and unambiguous, and thus easy to implement and follow. Essentially, UVAD is transductive learning, i.e., model training and testing are relatively independent. Numerous studies have shown that the performance of models in the normality learning phase on the agent task does not show a positive correlation with downstream anomaly detection. Due to the diversity of events, the spatial-temporal features of normal and abnormal samples overlap, and UVAD usually fails to actively recognize discriminative features and incorrectly learns some shared patterns when only regular events are available for model training (Liu et al., 2023f). In addition, Park et al. (Park et al., 2020) pointed out that overpowered deep neural networks may be able to effectively reason about unseen anomalous events during the testing phase due to overgeneralization performance, which may lead to underdetection. In response, the researchers proposed a memory network enhancement approach to weaken the model’s ability to generalize representations of anomalies by recording prototypical features.
@@
WsVAD uses weakly semantic video-level labels to supervise the output of strong semantic frame-level labels, i.e., frame-by-frame anomaly scores, by the sequencing model, thus enabling temporal localization of anomalous events (Liu et al., 2022d; Karim et al., 2024). The first weakly-supervised approach is the multiple instance ranking framework proposed by Sultani et al. (Sultani et al., 2018) in 2018, which lays out the basic MIL architecture of the WsVAD route, and whose concurrently publicly available UCF-Crime dataset has become the most widely used weakly-supervised benchmark. They consider a video as a collection of multiple examples (video clips), where clips containing abnormal frames are positive examples, while clips with all normal frames are labeled as negative examples. Obviously, a normal video with label 0 produces all examples called negative bags, and the example-level labels are all 00. An abnormal video with the label $1$ constitutes a positive packet, which contains both positive and negative instances. Inspired by multiple instance learning, WsVAD aims to train a scoring model to output the anomaly scores of each example using video-level labels. The authors introduce a MIL ranking loss inspired by the hinge loss, which encourages the model to output high anomaly scores close to 1 for anomalous clips by maximizing the difference between the anomaly scores of the largest-scoring instances in the positive and negative bags while scoring regular clips as close to 0 as possible. In fact, WsVAD does not belong to any of the classes of out-of-distribution detection tasks introduced in Section 2.1, but is rather a type of multiple instance learning under weak semantic labeling supervision.
@@
Compared to UVAD, the weakly supervised approach introduces anomalous videos in the training set and provides video-level labels for all training samples. Although anomalous examples are diverse and unenumerable in the real world, noteworthy anomalous events in specific scenarios are usually limited and easy to collect, such as thefts, robberies, and traffic accidents, which can be obtained in large quantities from surveillance IoT systems and online video platforms. In addition, the labor cost of video-level tagging, which only requires marking whether a video contains an anomaly without worrying about the specific timing location (frame-level labeling) and detailed anomaly category, is usually affordable. For example, the UCF-Crime dataset is much larger than the one used by UVAD, but only 1,900 discrete $[0,1]$ labels need to be provided. Due to the introduction of an additional human prior and the fact that the model has seen anomalous events during the training phase, WsVAD results are typically more reliable than UVAD, achieving excellent and consistent performance in detecting specific anomalous events. As a result, WsVAD has become a mainstream VAD scheme and is considered as the most promising research route for deployment in intelligent surveillance systems. The latest research attempts to mine anomaly-related clues from audio or subtitle text accompanying video frames, and proposes the multimodal WsVAD (Yang et al., 2024b).
@@
FuVAD attempts to learn anomaly classifiers directly from large-scale raw videos without any editing and labeling (Pang et al., 2020; Zaheer et al., 2022). Specifically, FuVAD’s training data contains both positive and negative samples, and due to the low frequency of anomalous examples compared to regular events, the anomalous samples to be detected can be regarded as outliers with different patterns from the main data. In essence, the FuVAD model is transductive learning and does not follow the training-testing process of UVAD. Fully unsupervised methods have become a research hotspot in the internet era, where data preparation is costly, by virtue of the fact that they do not require any constraints on the training data, and can be used to train the model by directly accessing huge amounts of videos from the real world (Liu et al., 2024h).
@@ 2.4. Emerging Research Tasks
2.5. VAD with Multimodal Large Language Models
@@
Large Language Models (LLMs) like Generative Pre-Training (GPT) (Radford, 2018) exhibit outstanding zero-shot learning and multimodal information processing abilities, showing great potential in VAD. Research has shown that multimodal LLMs can learn prototype patterns of normal events without training and describe any anomalies in open-set settings, significantly improving the generality and adaptability of VAD models. Zanella et al. (Zanella et al., 2024) proposed Language-based VAD (LAVAD), an unsupervised learning paradigm that leverages pre-trained LLMs and existing VLMs to train video anomaly detectors. They used VLMs to generate textual descriptions of video frames and designed a prompt mechanism to unlock LLMs’ potential in temporal aggregation and anomaly score estimation, enabling direct VAD execution. Lv et al. (Lv and Sun, 2024) introduced Video-LLaMA into VAD, aiming to break threshold limitations and improve model interpretability. The authors proposed a three-stage training method to improve the training efficiency of VLLM. In the AnomalyRuler, during the inductive stage, a small set of normal reference videos was provided to the LLM, enabling it to summarize normal patterns to induce rules for anomaly detection. In the Deduction stage, these induced rules were applied to detect anomalous frames in test videos. Hawk in (Tang et al., 2024) uses an interactive VLM to accurately interpret video anomalies, answering VAD-related questions. It constructs an auxiliary consistency loss within the motion and video space, guiding the video branch to focus on motion modalities and establishing explicit supervision between actions and language to improve the accuracy of interpretation. Zhang et al. (Zhang et al., 2024c) developed a large-scale multimodal VAD instruction tuning benchmark called VAD-Instruct50k, used to build unbiased and interpretable VAD systems, as well as Holmes-VAD for anomaly event localization and interpretation.